# Measuring Compliance of Consent Revocation on the Web

### v2

@@ 4.1.2 Analyzing interfaces for consent revocation (RQ1)
(4) After rejection: To compare the behavior of the website after initial rejection (opt-out) with that after the consent is explicitly revoked, we clean the browser and record the options to reject consent on the website. If a consent banner is present, we reject all cookies.
@@
Amid all these steps, we also take screenshots, in an automated manner, at different stages by scrolling to the bottom of the pages or navigating to the settings in order to get the important information like keywords leading to revocation. Using these inputs and screenshots, we categorize the websites based on the revocation options, where such options are provided, and the number of steps required to revoke consent. We then map these categories to the legal requirements for GDPR and consent revocation (from §3) to identify and measure potential violations of the law.
4.1.3 Effect of revocation on AA cookies (RQ2)
@@
In background, we collect cookies stored in the browser at each of the four stages. Since only some cookies require consent (see §3), we followed Bouhoula et al. [68] and consider that only cookies classified as Advertising or Analytics by CookieBlock [5] require user consent. If we detect AA cookies upon initial landing, after rejection or after revocation, we map it to legal requirements and identify a potential violation with the law. Similarly, we analyse whether the number of AA cookies increases after revocation w.r.t. other stages of our analysis.
4.2 Datasets for RQ3 and RQ4
@@
In websites that allowed revocation, different methods were used to store the modified consent and communicate it over the network. To detect the presence of CMPs that implement the IAB TCF [34], and therefore use a standardised format for storing consent (Transparency and Consent String or TCString), prior works [4, 40] used the __tcfapi function provided by TCF. We, therefore, performed an automated preliminary study to test this detection method. We also identify the detection methods for the use of OneTrust, which is the most popular CMP on Tranco top-20k websites [69].
4.2.1 Preliminary study
@@
We performed an automated crawl querying for __tcfapi, a function that must be provided to use TCF’s functionalities, on the top 200 websites in the EU region, according to Tranco list [65]. Out of 161 websites that were reachable, we found __tcfapi on only 32 (20%) of websites [^2] Similarly to prior work [4]. , we found that __tcfLocator is present on only 1 extra website where the __tcfapi is not present. We therefore do not test for this method in this work.. We therefore further explored other ways to detect CMPs by analysing the implementations of the most popular CMP [69], OneTrust. By examining websites with OneTrust and its documentation [70], we found that this CMP (1) stores user consent in a very specific format in a cookie named OptanonConsent, and (2) maintains a JavaScript variable, OneTrustActiveGroups (OTAG), which can be queried to get the current consent string. However, the format of the consent string accessible through these methods does not respect the official IAB Europe TCF TCString format.
4.2.2 Website selection
@@
We use the methods developed in the preliminary study to detect websites that use TCF and OneTrust from top-200 websites, resulting in dataset DS2 that contained only 56 websites with CMPs. We randomly chose additional 1000 websites between the rank 200 and 5000, and detected 225 websites that implemented either TCF or OneTrust, resulting in a set of 281 websites (DS2 and DS3).
@@
Figure 2: Data collection pipeline: to address RQ3 (validity and consistency of consent), the collected sets of consent strings are highlighted in orange, while for RQ4 (access and communication of consent to 3rd-parties), the sets of scripts parties involved in the communication of consent are highlighted in violet. The objects in green represent our methods for collecting the data, and the boxes in yellow are further described in Table III.
4.2.3 Data collection
@@
Figure 2 shows the data collection pipeline for RQ3 and RQ4 that we describe next. The data is collected during all four stages of the experiment (highlighted as “Save background data” in Figure 1). We collect the consent strings (both TCString and consent variable specific to OneTrust) for further analysis of validity and consistency of consent storage (RQ3), and complement it with the information about scripts that access consent strings via APIs, scripts that initiate the requests and third-parties with whom consent is shared for the analysis of communication of consent (RQ4).
@@
4.2.4 Collecting stored consent strings and third-parties that access them
@@
The current TCF v2.2, does not define the specific location where the user consent should be stored. Therefore, we collect all cookies and localStorage objects’ values that match the TCString format. Additionally, we collect the OptanonConsent cookie that OneTrust uses to store a user’s consent. We also inject scripts in the webpage to query __tcfapi and OTAG to analyse consent strings that third-parties would receive if they query these APIs, and examine network logs to detect and store TCStrings shared in the requests and responses. All the collected consent strings are visualised in Figure 2 in orange color.
@@
Third-parties can access consent strings via APIs or receive them via network requests. We capture (1) third-party scripts that fetch consent string via __tcfapi or OTAG; (2) HTTP requests that contain consent strings in TCString format (we do not capture OneTrust-specific strings in the traffic because they do not have a well defined format). We store all initiators and receivers of consent strings (see violet stores in Figure 2).
@@
4.2.5 Detecting and extracting TCStrings in network logs
@@
To detect the TCString in network logs, we look at different parts of the request/response pair. Previous works have extracted TCStrings only from specific URL parameters, such as gdpr_consent [4] following TCF v1.1. A follow-up work [40] checked for outgoing TCString in all URL parameters of HTTP requests and also in the HTTP cookie headers, since TCF v2.2 does not specify URL parameters to be used. When parsing URLs, we extract all query parameters, for example for a URL https://www.site.com/zzz.jpg?ISBN=XXX&UID=ABC123, we first extract the values XXX and ABC123 and then check whether they correspond to the TCString format, according to TCF specification [71]. The orange boxes in Figure 2 represent the functions we used to extract TCStrings from different parts of the network logs that are further explained in Table III.
@@
Since TCStrings do not have any protection mechanisms, any third-party that receives a TCString, can potentially modify it and return a different TCString back to the browser. To detect inconsistencies in TCString between requests and responses, we analyse postData in POST requests, and the response data (HTML and JSON formats) from the servers. In each case, we record the request initiator to identify scripts that sent wrong consent strings. To analyse HTTP GET requests, we search for TCString in the URL parameters, while for POST methods, we extract the TCString from JSON object if the data is sent in this format. To analyse HTTP responses and detect when TCString is modified by the server, we analyse redirected URLs: if response type is JSON, we parse it; otherwise, we parse the returned HTML file to extract TCString from the script, image and iFrame elements. Our approach uncovers more instances of TCStrings as compared to previous works [4, 40], since we detect 4,445 TCStrings in URL parameters, and an additional 3,423 TCStrings in the postData, 52 TCStrings in JSON objects and 170 TCStrings in HTML files. We also record requests where the cookie containing TCStrings is sent to third-parties.
Function in Fig. 2 Extracting TCString
@@
parseURL (URL) Extract all query parameters from URL and keep all parameter values that match a TCString format.
@@
parseJSON (obj) Extract all {key:value} pairs present in obj and keep all values that match a TCString format.
@@
parseHTML (file) Extract URLs for <img>, <iframe> and <script> in file and call parseURL to extract TCStrings.
TABLE III: Parsing the network logs for extracting TCStrings
4.2.6 Wrappers for __tcfapi and OTAG
@@
To observe who is accessing the __tcfapi, we override the function to record all the scripts which accessed it. We also record what command was requested through the call, e.g., “ping”, “getTCData” or “getEventHandler”. A similar approach was adopted by Matte et al. [4], however, they observed calls made to the APIs in an older version of TCF and their analysis did not include OneTrust. We override getter and setter methods of the OTAG variable to record which domain is writing/updating it (mostly by a designated CMP script) and reading/accessing it.
4.2.7 Communicating consent modification
@@
We check the network logs to investigate if third-parties that received the TCString after acceptance had also received the updated consent after revocation. We do this by observing requests to third-party URLs, and if the request contained the TCString either in the URL or in the POST data as described in the Figure 2.
4.2.8 Identifying responsible parties
@@
We record the initiators of all the network requests, and use this information to identify which third-party script changed the consent. We match script’s domain of provenance with the CMP name decoded from the TCStrings. If the name and domain do not match, in case of inconsistencies, we manually check the domains for cases where the CMP names are either shortened or have words like ‘privacy center’ or ‘CDN’. If we do not identify any such indicator that the third-party could be the CMP, we classify the responsible party just as a third-party.
4.3 Limitations
4.3.1 Accuracy of AA cookie classification
@@
Even though CookieBlock [5] has better accuracy than other tools for classifying cookie purposes, it might not be entirely accurate: a website may declare some of the cookies, classified as AA, as necessary in their cookie policies or banners. Such cookies, classified by CookieBlock, will be, nevertheless, labeled as AA and considered to require consent.
@@
4.3.2 Not capturing network logs while searching for revocation option on secondary pages
@@
When reaching the revocation settings takes multiple steps, for a few seconds, log-entries with the old consent (after acceptance) are also saved along with the network logs meant for “after revocation” stage. We remove these log-entries by searching for inconsistencies after the consent is revoked. However, the set of log-entries misclassified as “after revocation” are not checked for inconsistencies and we may have missed a few cases similar to the ones mentioned in Table VIII in the Appendix.
@@
4.3.3 Unable to capture access to consent when the third-parties use event listeners
@@
Callback functions returned as an event listener by __tcfapi does not have a standard format or name for us to track the access requests made to the API. Similarly, OneTrust suggests third-party ad-systems like Google Tag Manager to implement their own event listeners to be updated when the user changes the consent. Hence, we do not track API accesses when event listeners are used.
4.3.4 Determining description of purposes in OneTrust
@@
We do not know how the purposes displayed by OneTrust or any CMP used by the website are mapped to the user’s actual consent. Since there is no particular format used to store the consent string, it can vary between different CMPs, and within websites using the same CMP, as well.
4.3.5 Delay in registering user’s consent
@@
While collecting network logs to assess the communication of consent revocation to third parties, we observed (see Section .1) that some websites may take time to register that a user revoked their initial consent. To account for this, we exclude cases with delays in updating consent strings from our list of violations. However, we include cases where some third parties correctly include the updated consent string in their network requests, while others still send the old consent (positive consent string). Because the implementation for notifying third parties is server-side, this may result in some false positive violations on our end.
5 Revocation Interface and its Compliance
@@
This section examines user interfaces for revoking consent to determine if they potentially violate EU legal requirements, addressing RQ1. We first categorize the $161161$ reachable websites out of top 200 (datasets DS1 and DS2 from Table II), based on their consent banners: 108 (67%) websites display a consent banner to the users on their first visit, 8 (5%) websites displayed a banner with no option, not allowing neither to accept nor reject consent, and 45 (28%) websites did not display any banner. [^3] This can be explained by the fact that these websites were hosted outside of the EU and did not yet implement compliance solutions with EU laws. We further manually classify 161 websites based on the interface provided to revoke consent: 120 (74.5%) out of 161 websites provide users option to revoke consent within the same interface, medium or website where the consent request took place, or navigates to related pages for revoking consent; 32 (19.8%) websites provided users options to withdraw consent via different interface, where the options to revoke consent are present outside the interface or medium from where the consent request took place, and 9 (5.6%) websites offer no revocation option.
@@
Consent banner 108 9 63 6 21 8 0 1 0
No option banner 8 0 0 0 5 2 0 1 0
Without banner 45 0 4 5 7 17 3 0 4 [+5]
Total 161 9 67 11 33 27 3 2 9
@@
TABLE IV: Prevalence and type of consent revocation options on 161 websites (rank 1-200) from DS1. We use different colors to represent the level of compliance of the detected practices. Green color represents zero or one steps, yellow represents two or more steps, and pink and red represent violations of varying severity.
@@
Table IV summarizes the results for these three categories. To further evaluate legal compliance, for websites that provide revocation within the same interface, we count the number of steps required to revoke consent based on Figure 1. Figure 6 (in the Appendix) shows examples of these categories.
5.1 Results
5.1.1 Compliant revocation interface implementation (zero or one steps)

### v3

@@ 4.1.2. Analyzing interfaces for consent revocation (RQ1)
(4) After rejection: To compare the behavior of the website after initial rejection (opt-out) with that after the consent is explicitly revoked, we clean the browser (includes clearing history, cookies, cache, passwords, form data, and site settings to ensure consistency in each measurement) and record the options to reject consent on the website. If a consent banner is present, we reject all cookies.
@@
Amid all these steps, we also take screenshots in an automated manner at different stages by scrolling to the bottom of the pages or navigating to the settings in order to get the important information like keywords leading to revocation. Using these inputs and screenshots, we categorize the websites based on the revocation options, where such options are provided, and the number of steps required to revoke consent. We then map these categories to the legal requirements for GDPR and consent revocation (from §3) to identify and measure potential violations of the law. To reduce manual bias in our analysis, labeling was conducted by one author (twice, three months apart, in March and June of 2024.), with category examples co-developed with a legal scholar. Several examples were jointly reviewed by both a legal expert and a computer scientist.
4.1.3. Effect of revocation on AA cookies (RQ2)
@@
In the background, we collect cookies stored in the browser at each of the four stages. Since only some cookies require consent (see §3), we followed Bouhoula et al. (Bouhoula et al., 2024) and consider that only cookies classified as Advertising or Analytics (AA) by CookieBlock (Bollinger et al., 2022) require user consent. If we detect AA cookies upon initial landing, after rejection, or after revocation, we map it to legal requirements and identify a potential violation of the law. Similarly, we analyse whether the number of AA cookies increases after revocation w.r.t. other stages of our analysis.
In this analysis, we do not include the websites without a banner because we need to compare cookies after explicit acceptance and after revocation. Websites without a banner do not allow a user to accept consent, and therefore, modifying consent on such websites would not mean “revocation” in legal terms.
4.2. Datasets for RQ3 and RQ4
@@
In websites that allowed revocation, different methods were used to store the modified consent and communicate it over the network. To detect the presence of CMPs that implement the IAB TCF (Europe, 2025), and therefore use a standardised format for storing consent (Transparency and Consent String or TCString), prior works (Matte et al., 2020a; Zhang et al., 2024) used the __tcfapi function provided by TCF. We, therefore, performed an automated preliminary study to test this detection method. We also identify the detection methods for the use of OneTrust, which is the most popular CMP on Tranco top-20k websites (Hils et al., 2021).
4.2.1. Preliminary study
@@
We performed an automated crawl querying for __tcfapi, a function that must be provided to use TCF’s functionalities, on the top-200 domains (Tranco, 2023). Out of 158 websites that were reachable, we found __tcfapi on only 32 (20%) of websites [^1] Similar to prior work (Matte et al., 2020a), we found that __tcfLocator is present on only 1 website where __tcfapi is not present. We therefore do not test for this method here.. We therefore further explored other ways to detect CMPs by analysing the implementations of the most popular CMP (Hils et al., 2021), OneTrust. By examining websites with OneTrust and its documentation (OneTrust, [n. d.]), we found that this CMP (1) stores user consent in a very specific format in a cookie named OptanonConsent, and (2) maintains a JavaScript variable, OneTrustActiveGroups (OTAG), which can be queried to get the current consent string. However, the format of the consent string accessible through these methods does not respect the official IAB Europe TCF TCString format.
4.2.2. Website selection
@@
We use the methods developed in the preliminary study to detect websites that use TCF and OneTrust from top-200 websites, resulting in dataset DS2 with only 56 websites with CMPs. We randomly chose an additional 1000 websites between the rank 200 and 5000 and detected 225 websites with either TCF or OneTrust resulting in a set of 281 websites (DS2 and DS3).
@@
Figure 2. Data collection pipeline: to address RQ3 (validity and consistency of consent), the collected sets of consent strings are highlighted in orange, while for RQ4 (access and communication of consent to 3rd-parties), the sets of scripts parties involved in the communication of consent are highlighted in violet. The objects in green represent our methods for collecting the data, and the boxes in yellow are further described in Table 3.
4.2.3. Data collection
@@
Figure 2 shows the data collection pipeline for RQ3 and RQ4. The data is collected during all four stages of the experiment (highlighted as “Save background data” in Figure 1). We collect the consent strings (both TCString and consent variable specific to OneTrust) for further analysis of validity and consistency of consent storage (RQ3), and complement it with the information about scripts that access consent strings via APIs, scripts that initiate the requests and third-parties with whom consent is shared for the analysis of communication of consent (RQ4).
@@
4.2.4. Collecting stored consent strings and third-parties that access them
@@
The current TCF v2.2, does not define the specific location where the user consent should be stored. Therefore, we collect all cookies and localStorage objects’ values that match the TCString format. Additionally, we collect the OptanonConsent cookie that OneTrust uses to store a user’s consent. We also inject scripts in the webpage to query __tcfapi and OTAG to analyse consent strings that third-parties would receive if they query these APIs and examine network logs to detect and store TCStrings shared in the requests and responses. All the collected consent strings are visualised in Figure 2 in orange color.
@@
Third-parties can access consent strings via APIs or receive them via network requests. We capture (1) third-party scripts that fetch consent string via __tcfapi or OTAG; (2) HTTP requests that contain consent strings in TCString format or OTAG. However, some requests might not have been recorded since they do not have a well-defined format. We store all initiators and receivers of consent strings (see violet stores in Figure 2).
@@
4.2.5. Detecting and extracting TCStrings and specific categories from OneTrust in network logs
@@
To detect the TCString in network logs, we look at different parts of the request/response pair. Previous works have extracted TCStrings only from specific URL parameters , such as gdpr_consent (Matte et al., 2020a) following TCF v1.1. A follow-up work (Zhang et al., 2024) checked for outgoing TCString in all URL parameters of HTTP requests and also in the HTTP cookie headers since TCF v2.2 does not specify URL parameters to be used.
When parsing URLs, we extract all query parameters, for example for a URL https://www.site.com /zzz.jpg?ISBN=XXX&UID=ABC123, we first extract the values XXX and ABC123 and then check whether they correspond to the TCString format, according to TCF specification (IAB, 2024). The orange boxes in Figure 2 represent the functions we used to extract TCStrings from different parts of the network logs that are further explained in Table 3.
@@
Since TCStrings do not have any protection mechanisms, any third-party that receives a TCString, can potentially modify it and return a different TCString back to the browser. To detect inconsistencies in TCString between requests and responses, we analyse postData in POST requests, and the response data (HTML and JSON formats) from the servers. In each case, we record the request initiator to identify scripts that sent the wrong consent strings. To analyse HTTP GET requests, we search for TCString in the URL parameters, while for POST methods, we extract the TCString from JSON object if the data is sent in this format. To analyse HTTP responses and detect when TCString is modified by the server, we analyse redirected URLs: if response type is JSON, we parse it; otherwise, we parse the returned HTML file to extract TCString from the script, image and iFrame elements.
Our approach uncovers more instances of TCStrings as compared to previous works (Matte et al., 2020a; Zhang et al., 2024) — we detect 4,445 TCStrings in URL parameters, 3,423 TCStrings in the postData, 52 TCStrings in JSON objects and 170 TCStrings in HTML files. We also record requests where the cookie containing TCStrings is sent to third-parties.
Func. in Fig. 2 Extracting TCString
@@
parseURL (URL) Extract all query parameters from URL and keep all values that match a TCString format or contain OneTrust cookie categories.
@@
parseJSON (obj) Extract all {key:value} pairs present in obj and keep all values that match a TCString format or contain OneTrust cookie categories.
@@
parseHTML (file) Extract URLs for img, iframe and script in file and parse them to extract TCStrings or OneTrust cookie categories.
Table 3. Extracting TCStrings and OneTrust consent string
4.2.6. OneTrust-specific encoding of consent
Differently from other CMPs, OneTrust CMP does not use the TCString but uses a specific format to store and communicate consent: the consent can be stored as key-value pairs in two locations, either in the OptanonConsent cookie or in the OTAG variable. The cookie values contain encoded categories of cookies (e.g., “analytics” or “advertisement”), that follow the format such as “X:1;Y:0;Z:1”, where X, Y and Z indicate cookie categories and 0 or 1 indicate rejection or acceptance to use the specific category in the consent. The values of the OTAG list all the allowed cookie categories (called “active groups” in OneTrust documentation); however, the encoded format of such categories is not specified in OneTrust documentation. Our analysis of inconsistency between the OptanonConsent cookie value and OTAG variable (section 7) revealed that when two values are not consistent, OptanonConsent cookie value contained correct record of user consent. We therefore opted for extracting the allowed categories from the OptanonConsent cookies and search all such encoding (only X and Z in the example above) in the URL query parameters to record the consent strings being sent to third-parties.
However, the variable in the query parameter having the consent string is not consistent. This leads to false positives where the active groups are represented using numbers like ‘1’, ‘2’, and ‘3’ or “1:1,2:0,3:0,4:1". These representations are hard to distinguish from query parameters used for purposes other than sharing consent. We manually analyzed such websites, and hence, could have missed some websites due to human error.
4.2.7. Wrappers for __tcfapi and OTAG
@@
To identify __tcfapi access, we override the function to record all the scripts that access it. We also record what command was requested, e.g., “ping”, “getTCData” or “getEventHandler”. A similar approach was adopted by Matte et al. (Matte et al., 2020a). However, they observed calls made to the APIs in an older version of TCF, and their analysis did not include OneTrust. We override getter and setter methods of the OTAG variable to record which domain is writing/updating it (mostly by a designated CMP script) and reading/accessing it.
4.2.8. Communicating consent modification
@@
We check the network logs to investigate if third-parties that received the TCStringand One Trust consent string (containing allowed cookie categories) after acceptance had also received the updated consent after revocation. We do this by observing requests to third-party URLs and if the request contained the TCString or One Trust consent string either in the URL or in the POST data as described in Figure 2.
4.2.9. Identifying responsible parties
@@
We record the initiators of all the network requests and use this information to identify which third-party script changed the consent. We match the script’s domain of provenance with the CMP name decoded from the TCStrings. If the name and domain do not match, in case of inconsistencies, we manually check the domains for cases where the CMP names are either shortened or have words like ‘privacy center’ or ‘CDN’. If we do not identify any such indicator that the third-party could be the CMP, we classify the responsible party just as a third-party.
4.3. Limitations
4.3.1. Accuracy of AA cookie classification
@@
Even though CookieBlock (Bollinger et al., 2022) has better accuracy than other tools for classifying cookie purposes, it might not be entirely accurate: a website may declare some of the cookies, classified as AA, as necessary in their cookie policies. Such cookies, classified by CookieBlock, will be, nevertheless, labeled as AA and considered to require consent.
@@
4.3.2. Not capturing network logs while searching for revocation option on secondary pages
@@
When reaching the revocation settings takes multiple steps, for a few seconds, log-entries with the old consent (after acceptance) are also saved along with the network logs meant for “after revocation” stage. We remove these log-entries by searching for inconsistencies after the consent is revoked. However, the set of log-entries misclassified as “after revocation” are not checked for inconsistencies, and we may have missed a few cases similar to the ones mentioned in Table 8 in the Appendix.
@@
4.3.3. Unable to capture access to consent when the third-parties use event listeners
@@
Callback functions returned as an event listener by __tcfapi do not have a standard format or name to track the access requests made to the API. Hence, we do not track API accesses when event listeners are used.
4.3.4. Determining description of purposes in OneTrust
@@
We do not know how the purposes displayed by OneTrust or any CMP used by the website are mapped to the user’s actual consent. Since there is no particular format used to store the consent string, it can vary between different CMPs and within websites using the same CMP.
4.3.5. Delay in registering user’s consent
@@
While collecting network logs to assess the communication of consent revocation to third parties, we observed (see Section A) that some websites may take time to register that a user revoked their initial consent. To account for this, we exclude cases with delays in updating consent strings from our list of violations. However, we include cases where some third parties correctly include the updated consent string in their network requests, while others still send the old consent (positive consent string). As the third-party notification happens server-side, this may result in some false-positives on our end.
4.4. Data Availability and Disclosure
As described in §4.1 and Figure 1, we collect background data, screenshots, and website labels. The complete dataset, including the network logs and the crawler, is provided in the supplementary material (Anonymous, 2025). Some examples are shown in the Appendix.
As of the date of submission of this paper, 28 February 2025, we have notified all the companies who own the domains that we explicitly mention in the main text of our paper. We will add information on the responses we receive, in the final draft.
5. Revocation Interface and its Compliance
@@
To address RQ1, we examined user interfaces for revoking consent and checked if they (potentially) violated EU legal requirements. We first categorize the $158158$ reachable websites (datasets DS1 and DS2 from Table 2), based on their consent banners: 108 (67%) websites display a consent banner to the users on their first visit, 8 (5%) websites displayed a banner with no option, not allowing to accept or reject consent, and 45 (28%) websites did not display any banner. [^2] This can be explained by the fact that these websites were hosted outside of the EU and did not yet implement compliance solutions with EU laws. We manually classify the 158 websites based on the interface to revoke consent: 120 (74.5%) websites provide users option to revoke consent within the same interface where the consent request took place, or navigates to related pages for revoking consent; 32 (19.8%) websites provided options via different interface, where such options are present outside the interface or medium, where the consent request took place, and 9 (5.6%) websites offer no revocation option.
@@
Consent banner 105 8 61 4 23 8 1 0 0
No option banner 8 0 0 0 5 3 0 0 0
Without banner 45 0 7 2 7 17 0 3 4 [+5]
Total 158 8 68 6 35 28 1 3 9
@@
Table 4. Prevalence and type of consent revocation options on 158 websites (rank 1-200) from DS1. We use different colors to represent the level of compliance of the detected practices. Green color represents zero or one steps, yellow represents two or more steps, and pink and red represent violations of varying severity.
@@
Table 4 summarizes the results for these three categories. To further evaluate legal compliance, for websites that provide revocation within the same interface, we count the number of steps required to revoke consent based on Figure 1. Figure 6 (in the Appendix) shows example screenshots for each of these categories.
5.1. Results
5.1.1. Compliant revocation interface (zero or one steps)